# Post-detection inference for sequential changepoint localization

### v1

@@ 15            $\mathcal{C}=\mathcal{C}\cup\{t\}$
Step 1 [Establish the relationship between the level of testing ${H}_{0,T}$ and the conditional coverage]: For each $t$, define $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$. Then, the level of our test for ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Now,
$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$ $=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
@@
$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{\infty}(\tau<T)$
$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T).$
The inequality in the third line holds because
$\mathbb{P}_{T}(T\in\{t\geq 1:M_{t}\leq C_{t}^{N}\},\tau<T)\leq\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$
where the last equality holds since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$ and by the same argument, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{\mathbb{P}}_{\infty}(\tau\geq t)$ for all $t$. So, we obtain
@@
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$ (2.16)
Then, it follows that
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$ (2.17)
@@
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t)=1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ (2.18)
Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$ and hence, $C_{t,L}\geq C_{t,\infty}$.
Thus, we obtain $\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$, for all $t$.
Therefore, from (2.15), we have $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
@@ 3 Known pre-change but unknown post-change distribution
Notation.
For $t\in\mathbb{N}$, let $\mathbb{P}_{\theta,t,\theta^{\prime}}$ denote the joint distribution of the sequence of independent random variables $\{X_{n}\}_{n\in\mathbb{N}}$, when $X_{1},\dots,X_{t-1}$ are from $F_{\theta}$ (with density $f_{\theta}$) and $X_{t},X_{t+1},\dots$ are from $F_{\theta^{\prime}}$ (with density $f_{\theta^{\prime}}$). As corner cases, let $\mathbb{P}_{1,\theta^{\prime}}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta^{\prime}}$ and $\mathbb{P}_{\theta,\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta}$. Let $\mathbb{E}_{\theta,t,\theta^{\prime}}$ denote the expectation under $\mathbb{P}_{\theta,t,\theta^{\prime}}$, for $t\in\mathbb{N}\cup\{\infty\}$, for $t\in\mathbb{N}\cup\{\infty\}$ and $\mathbb{E}_{\theta,\infty}$ denote the expectation under $\mathbb{P}_{\theta,\infty}$.
@@
Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$. However, note that the sample size used for estimate might be too small to have a reliable estimate. To address this issue, we would use interval estimates instead of point estimates so that we can hope to achieve a predetermined control over the error. Notably, conventional confidence intervals are invalid in this context, since $\tau$ depends on the data. Therefore, we rely on confidence sequences (Darling and Robbins,, 1967; Jennison and Turnbull,, 1989; Howard et al.,, 2021). A confidence sequence having coverage $1-c$ is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform coverage guarantee: $\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage. Otherwise, we construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage. Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS has coverage probability $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 3.1.
@@
where $r_{t}:=\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N$. If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$, we can have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
@@
Note that, unless $\mathcal{S}^{\prime}_{\tau-t+1}$ is a finite set, it can be computationally challenging to approximate the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$. However, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient. We provide details in Section 3.2 and a concrete Gaussian mean shift example in Section 3.3.
@@ 22            $\mathcal{C}=\mathcal{C}\cup\{t\}$
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\delta\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{1-\delta},\text{ when }\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
@@ 3.1 Prototypical choice of test statistic $M_{t}$
We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$. The point estimate of $T$ can be naturally extended to the following form:
@@
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{ if }t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (3.7)
@@ 3.2 Computationally and statistically efficient subroutines
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$.
@@
$V_{t,L}^{j}:=\begin{cases}-\infty&\text{ if }t>t_{2}^{j}\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{ if }t\leq t_{2}^{j}\leq L\\ \infty&\text{ if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}$ (3.9)
@@
$V_{t,\infty}^{j}:=\begin{cases}\displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{ if }t\leq t_{2}^{j}\text{ and }t_{1}^{j}<\infty\\ -\infty&\text{ if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \end{cases}$ (3.10)
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, as we explain in Section 3.3). We define the confidence interval as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (3.11)
@@
$\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$ $\geq\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t);M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$

### v2

@@ 15            $\mathcal{C}=\mathcal{C}\cup\{t\}$
First note the fact that $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
Step 1 [Relate the level of testing ${H}_{0,T}$ and the conditional coverage]: Defining $C_{t,L}=\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$, the level of testing ${H}_{0,T}$ is $P_{T}(M_{T}>C_{t,L}).$ Thus,
$\mathbb{P}_{T}(M_{T}\leq C_{t,L})$
$=\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:M_{t}\leq C_{t,L}\})$
@@
$\leq\mathbb{P}_{T}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq C_{t,L}\}\mid\tau\geq T)\mathbb{P}_{T}(\tau\geq T)+\mathbb{\mathbb{P}}_{T}(\tau<T)$
$=\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)+1-\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$
where we twice invoked $\mathbb{P}_{T}(\tau<T)=\mathbb{\mathbb{P}}_{\infty}(\tau<T)$ as explained earlier. So, we obtain
@@
$\mathbb{P}_{t}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c\text{ for all }t.$
Since $\hat{\alpha}_{t}$ is independent of $M_{t}$, it follows that
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t})\geq\frac{\lceil(1-\hat{\alpha}_{t})(B+1)\rceil}{B+1}\geq 1-\hat{\alpha}_{t},\text{ almost surelyfor all }t.$
@@
$\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})=\mathbb{E}_{t}(\mathbb{P}_{t}(M_{t}\leq C_{t,\infty}\mid\hat{\alpha}_{t}))\geq\mathbb{E}_{t}(1-\hat{\alpha}_{t})=1-\alpha\mathbb{P}_{t}(\tau\geq t).$
As noted earlier, $\mathbb{P}_{t}(\tau\geq t)=\mathbb{P}_{\infty}(\tau\geq t)$ and hence, $\mathbb{P}_{t}(M_{t}\leq C_{t,\infty})\geq 1-\alpha\mathbb{P}_{\infty}(\tau\geq t).$ Also, for all $L\in\mathbb{N}$, it follows from the definitions that $M_{t,L}^{j}\geq M_{t,\infty}^{j}$. Therefore, $C_{t,L}\geq C_{t,\infty}$. Thus, for all $t$,
$\mathbb{P}_{t}(M_{t}>C_{t,L})\leq\mathbb{P}_{t}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{\infty}(\tau\geq t)$ (2.16)
Therefore, combining (2.15) and (2.16), we obtain $\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
@@ 3 Known pre-change but unknown post-change distribution
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{\theta,t,\theta^{\prime}}$ denote the joint distribution of the sequence of independent random variables $\{X_{n}\}_{n\in\mathbb{N}}$, when $X_{1},\dots,X_{t-1}$ are from $F_{\theta}$ (with density $f_{\theta}$) and $X_{t},X_{t+1},\dots$ are from $F_{\theta^{\prime}}$ (with density $f_{\theta^{\prime}}$). As corner cases, let $\mathbb{P}_{1,\theta^{\prime}}$ denotes the joint distribution where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta^{\prime}}$ and $\mathbb{P}_{\theta,\infty}$ denotes the case where $X_{1},X_{2},\dots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{\theta}$. Let $\mathbb{E}_{\theta,t,\theta^{\prime}}$ denote the expectation under $\mathbb{P}_{\theta,t,\theta^{\prime}}$, for $t\in\mathbb{N}\cup\{\infty\}$, for $t\in\mathbb{N}\cup\{\infty\}$ and $\mathbb{E}_{\theta,\infty}$ denote the expectation under $\mathbb{P}_{\theta,\infty}$.
@@
Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$. However, note that the sample size used for estimate might be too small to have a reliable estimate. To address this issue, we would use interval estimates instead of point estimates so that we can hope to achieve a predetermined control over the error. Notably, conventional confidence intervals are invalid in this context, since $\tau$ depends on the data. Therefore, we rely on confidence sequences [4, 11, 9, 2]. A confidence sequence having coverage $1-c$ is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform coverage guarantee: $\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ If $\mathcal{A}$ controls PFA, we can simply construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta)$ for $\theta_{1}$ having $1-\beta$ coverage. Otherwise, we construct $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t})$, where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$, which have $1-\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$ coverage. Despite $r_{t}$ being random, it is based on independent simulations, and so it turns out that marginalizing over $r_{t}$ yields that the CS has coverage probability $\beta\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq t)$; we formalize this more explicitly in the proof of the Theorem 3.1.
@@
where $r_{t}:=\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. If $\mathcal{A}$ controls PFA at some known level $\delta\in(0,1)$, we can have the following confidence set, which is simpler (i.e., does not require $\tau_{i}$’s):
@@
We note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, one can approximate the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ by discretizing the parameter space into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum. Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient, and exact (i.e., does not require any approximation). We provide the details in Section 3.2. Additionally, we illustrate concrete examples of the Gaussian mean shift problem in Section 3.3 and the Poisson rate change problem in Appendix B.
@@ 22            $\mathcal{C}=\mathcal{C}\cup\{t\}$
$\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2})\geq 1-\alpha-\beta-\delta\leavevmode\nobreak\ \text{ , and }\leavevmode\nobreak\ \mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{2}\mid\tau\geq T)\geq 1-\frac{\alpha+\beta}{1-\delta},$
where the latter guarantee holds if $\mathbb{\mathbb{P}}_{\theta,\infty}(\tau\geq T)\neq 0.$
@@ 3.1 Prototypical choice of test statistic $M_{t}$
We define $\hat{\theta}_{1,j:n}$ to be a maximum likelihood estimator (MLE) of the post-change parameter $\theta_{1}\in\Theta_{1}$ based on data $X_{j},\cdots,X_{n}$. The point estimate of $T$ can be naturally extended to:
@@
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (3.7)
@@ 3.2 Computationally and statistically efficient subroutines
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$. In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Section B.1 for the Poisson distribution.
@@
$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (3.9)
@@
$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (3.10)
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section 3.3). We define the confidence set as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (3.11)
@@
$\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{V_{t,L}^{j}\}_{j=1}^{B})$
$\geq\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t,L}^{j}(\theta)\}_{j=1}^{B})$
$\geq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\frac{\alpha}{N}\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t);M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}).$
@@
In the next subsection, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with Poisson distribution in Appendix B.